# A Method for Handling Negative Similarities in Explainable Graph Spectral Clustering of Text Documents - Extended Version thanks: Supported by Polish Ministry of Science

### v1

@@ Abstract
Keywords:
Artificial Intelligence Machine Learning Graph spectral clustering document embeddings negative similarities
@@ 2 Previous work
where $S^{\prime}=S+I$ and $D^{\prime}=D+I$ ($I$ identity matrix).
@@ 4 TweetGloVe 0 0
Assume that we extend this space by the N+1st dimension, orthogonal to all other dimensions, and move the documents by a vector of length $b$ along this dimension. In Figure 1 it is illustrated as follows: Pick up two document embedding vectors, D1 and D2 (normalized to a unit length) and let the angle between them be $\alpha$ in the original space. Note that the three points D1, D2 and O (the origin of the coordinate system) span a hyperplane in the $N$ dimensional space, and this plane may be spanned by e.g. two mutually orthogonal unit vectors $\vec{OX1},\vec{OX2}$. We add the $(N+1)$st dimension and place the point $O^{\prime}$ on the $(N+1)$st axis at the distance $b$ from the origin. The vector $\vec{OO^{\prime}}$ is orthogonal to all vectors in the N-dimensional space, including $\vec{OD1},\vec{OD2}$. Let the angle between the vectors $\vec{O^{\prime}D1},\vec{O^{\prime}D2}$ representing documents in a “shifted space” be $\gamma`(b)$. Whereby (obviously) $\gamma(0)=\alpha$.
@@ 3.2 The problem of normalized Laplacians
Let us consider how serious is the risk of negative elements $D$. If all elements of $D$ were positive, the sum of all elements of $S$ would also be positive. So consider a bit abstract embedding into one-dimensional space. Then normalized document embedding vectors would be either [1] or [-1]. Let $p$ documents have the coordinate [1] and $q=n-p$ documents have the coordinate [-1]. In such a case $p(p-1)+q(q-1)$ ordered pairs of documents would have similarity 1, while $2pq$ would be of similarity -1. So, the sum of $S$ matrix would be $p(p-1)+q(q-1)-2pq=(p-q)^{2}-p-q=(p-q)^{2}-n$. In the worst case, the sum of $S$ would be $-n$. In higher dimensions, this would be less serious, but still, the risk of negativity remains, independently of the size of the document collection.
@@
Another solution would be adding a constant $c$ to each similarity ($s^{(pA)}_{ik}(c)=s_{ik}+c$. Again, no warranty that the ordering of all normalized similarities will be the same and hence that clustering result is the same.
@@
whereby $\max$ is computed over all off-diagonal elements of the $S$ matrix. $\cos$ is non-negative in the range $[0,\frac{\pi}{2}]$ while it is negative for greater angles. By dividing the actual angles between documents by the maximal angle, and multiplying with $\frac{\pi}{2}$ we scale all the angles into the non-negative cosine range. Now, the traditional normalized Laplacian is applicable. The ranking of similarities of combinatorial Laplacian is preserved completely, but again no warranty for the normalized similarities. The above formula can be generalized to:
@@
$c=1$ is for sure a reasonable choice, because arccos returns values in the range $[0,\pi]$ and dividing this result by 2 scales them into the required range $[0,\frac{\pi}{2}]$.

### v2

@@ 2 Previous work
where $S^{\prime}=S+I$ and $D^{\prime}=D+I$ (II - identity matrix).
@@ 3.1.1 Geometric Interpretation of $s_{ik}^{(pN)}$
Assume that we extend this space by the N+1st dimension, orthogonal to all other dimensions, and move the documents by a vector of length $b$ along this dimension. In Figure 1 it is illustrated as follows: Pick up two document embedding vectors, D1 and D2 (normalized to a unit length) and let the angle between them be $\alpha$ in the original space. Note that the three points D1, D2 and O (the origin of the coordinate system) span a hyperplane in the $N$ dimensional space, and this plane may be spanned by e.g. two mutually orthogonal unit vectors $\vec{OX1},\vec{OX2}$. We add the (N+1)(N+1)st dimension and place the point $O^{\prime}$ on the (N+1)(N+1)st axis at the distance $b$ from the origin. The vector $\vec{OO^{\prime}}$ is orthogonal to all vectors in the N-dimensional space, including $\vec{OD1},\vec{OD2}$. Let the angle between the vectors $\vec{O^{\prime}D1},\vec{O^{\prime}D2}$ representing documents in a “shifted space” be $\gamma`(b)$. Whereby (obviously) $\gamma(0)=\alpha$.
@@ 3.2 The problem of normalized Laplacians
Let us consider how serious is the risk of negative elements $D$. If all elements of $D$ were positive, the sum of all elements of $S$ would also be positive. So consider a bit abstract embedding into one-dimensional space. Then normalized document embedding vectors would be either [1] or [-1]. Let $p$ documents have the coordinate [1] and q=npq=n-p documents have the coordinate [-1]. In such a case p(p1)+q(q1)p(p-1)+q(q-1) ordered pairs of documents would have similarity 1, while $2pq$ would be of similarity -1. So, the sum of $S$ matrix would be $p(p-1)+q(q-1)-2pq=(p-q)^{2}-p-q=(p-q)^{2}-n$. In the worst case, the sum of $S$ would be n-n. In higher dimensions, this would be less serious, but still, the risk of negativity remains, independently of the size of the document collection.
@@
Another solution would be adding a constant $c$ to each similarity ($s^{(pA)}_{ik}(c)=s_{ik}+c)$. Again, no warranty that the ordering of all normalized similarities will be the same and hence that clustering result is the same.
@@
whereby $\max$ is computed over all off-diagonal elements of the $S$ matrix. $\cos$ is non-negative in the range $0,\frac{\pi}{2}]$ while it is negative for greater angles. By dividing the actual angles between documents by the maximal angle, and multiplying with $\frac{\pi}{2}$ we scale all the angles into the non-negative cosine range. Now, the traditional normalized Laplacian is applicable. The ranking of similarities of combinatorial Laplacian is preserved completely, but again no warranty for the normalized similarities. The above formula can be generalized to:
@@
$c=1$ is for sure a reasonable choice, because arccos returns values in the range $0,\pi]$ and dividing this result by 2 scales them into the required range $0,\frac{\pi}{2}]$.
@@
3.2.1 Transformation preserving some properties of normalized Laplacian
For combinatorial Laplacians, We mentioned a solution (14) $s_{ik}^{(pN)}=\frac{s_{ik}+c}{1+c}$ consisting in adding a constant $c$ to each similarity and then normalizing. While with normalized Laplacian, we do not have a warranty that the ordering of all normalized similarities will be the same and hence that clustering result is the same, this ordering can be preserved if the similarity matrix has reasonable properties mentioned below, so that related properties of $D^{-0.5}SD^{-0.5}$ will be preserved and the clustering shall not deviate significantly.
Let us investigate some details of the transformation $s_{ik}^{(pN)}$. Assume $s_{ik}$, as a product of unit document vectors $\mathbf{v}_{i},\mathbf{v}_{j}$, range from 0 to 1. Add a new dimension and make new unit document vector: $\mathbf{v}_{i}^{\prime}=(\mathbf{v}^{T}_{i},\sqrt{x})^{T}/\sqrt{1+x}$. New $s_{ik}^{\prime}=\frac{s_{ik}+x}{1+x}$. New $d_{ii}^{\prime}=\sum_{k}s_{ik}^{\prime}=\frac{d_{ii}}{1+x}+\frac{(n-1)x}{1+x}$
Define as a consequence the "normalized" $s_{ik}^{(v)}$:
$s_{ik}^{(v)}=\frac{s_{ik}^{\prime}}{\sqrt{d_{ii}^{\prime}}\sqrt{d_{kk}^{\prime}}}=\frac{\frac{s_{ik}+x}{1+x}}{\sqrt{\frac{d_{ii}}{1+x}+\frac{(n-1)x}{1+x}}\sqrt{\frac{d_{kk}}{1+x}+\frac{(n-1)x}{1+x}}}$ (20)
that is, by multiplication by $1+x$ the nominator and the denominator of the right-hand-side.
$s_{ik}^{(v)}=\frac{s_{ik}+x}{\sqrt{d_{ii}+(n-1)x}\sqrt{d_{kk}+(n-1)x}}$ (21)
Consider three documents, $i,k,\ell$. Let $i,k$ belong to the same cluster, and $\ell$ to the other cluster. It would be reasonable to assume therefore that the similarity between $i$ and $K$ is bigger than between $i$ and $\ell$, that is
$s_{ik}>s_{i\ell}$
Furthermore, the similarity between $i,k$ is reasonably bigger than average similarity between $k$ and anything else. Also, the similarity between $i,\ell$ is reasonably smaller than average similarity between $\ell$ and anything else.
Let us demonstrate that under these reasonable assumptions
$s_{ik}^{(v)}>s_{i\ell}^{(v)}$ (22)
for $x\geq 0$. The above expression is equivalent to:
$\frac{s_{ik}+x}{\sqrt{d_{ii}+(n-1)x}\sqrt{d_{kk}+(n-1)x}}>\frac{s_{i\ell}+x}{\sqrt{d_{ii}+(n-1)x}\sqrt{d_{\ell\ell}+(n-1)x}}$ (23)
The above expression is equivalent to:
$\frac{s_{ik}+x}{\sqrt{d_{kk}+(n-1)x}}>\frac{s_{i\ell}+x}{\sqrt{d_{\ell\ell}+(n-1)x}}$ (24)
Let us divide the denominator of both sides by $\sqrt{n-1}$
$\frac{s_{ik}+x}{\sqrt{\frac{d_{kk}}{n-1}+x}}>\frac{s_{i\ell}+x}{\sqrt{\frac{d_{\ell\ell}}{n-1}+x}}$ (25)
As we assumed already, $\frac{d_{kk}}{n-1}<s_{ik}$ that is $\frac{d_{kk}}{n-1}=s_{ik}-\delta_{k}$. Also $\frac{d_{\ell\ell}}{n-1}>s_{i\ell}$ that is $\frac{d_{\ell\ell}}{n-1}=s_{i\ell}+\delta_{\ell}$. Whereby $\delta_{k}>0,\delta_{\ell}>0$.
Hence the above is equivalent:
$\frac{s_{ik}+x}{\sqrt{s_{ik}-\delta_{k}+x}}>\frac{s_{i\ell}+x}{\sqrt{s_{i\ell}+\delta_{\ell}+x}}$ (26)
Divide the nominator and denominator of the left hand side by $s_{ik}+x$, and the nominator and denominator of the right hand side by $s_{i\ell}+x$
$\frac{1}{\sqrt{\frac{1}{s_{ik}+x}-\frac{\delta_{k}}{(s_{ik}+x)^{2}}}}>\frac{1}{\sqrt{\frac{1}{s_{i\ell}+x}+\frac{\delta_{\ell}}{(s_{i\ell}+x)^{2}}}}$ (27)
$\sqrt{\frac{1}{s_{i\ell}+x}+\frac{\delta_{\ell}}{(s_{i\ell}+x)^{2}}}>\sqrt{\frac{1}{s_{ik}+x}-\frac{\delta_{k}}{(s_{ik}+x)^{2}}}$ (28)
$\frac{1}{s_{i\ell}+x}+\frac{\delta_{\ell}}{(s_{i\ell}+x)^{2}}>\frac{1}{s_{ik}+x}-\frac{\delta_{k}}{(s_{ik}+x)^{2}}$ (29)
$\frac{\delta_{k}}{(s_{ik}+x)^{2}}+\frac{\delta_{\ell}}{(s_{i\ell}+x)^{2}}>\frac{1}{s_{ik}+x}-\frac{1}{s_{i\ell}+x}$ (30)
$\frac{\delta_{k}}{(s_{ik}+x)^{2}}+\frac{\delta_{\ell}}{(s_{i\ell}+x)^{2}}>\frac{(s_{i\ell}+x)-(s_{ik}+x)}{(s_{ik}+x)(s_{i\ell}+x)}$ (31)
$\frac{\delta_{k}}{(s_{ik}+x)^{2}}+\frac{\delta_{\ell}}{(s_{i\ell}+x)^{2}}>\frac{s_{i\ell}-s_{ik}}{(s_{ik}+x)(s_{i\ell}+x)}$ (32)
As $s_{i\ell}<s_{ik}$, the righthandside is smaller than zero. As
$\frac{\delta_{k}}{(s_{ik}+x)^{2}}+\frac{\delta_{\ell}}{(s_{i\ell}+x)^{2}}>0$ (33)
and
$0>\frac{s_{i\ell}-s_{ik}}{(s_{ik}+x)(s_{i\ell}+x)}$ (34)
we see that (32) is true, hence also (22).